a rectangular plate 2* 3 m is immersed in water such that its greatest and least depth are 6m and 4m respectively from the water surface . the total thrust on the plate is what?

Grade:12

2 Answers

Arun
5 years ago
Total pressure, F=w*a*y=9.81*650*6*2*3=22.9 N/cm2

Regards
Arun (askIITians forum expert)
Ashutosh Dangi
3 years ago
Given, size of the plate 2mx 3m and Greatest and least depths of the plate are 6m and 4m.
 
We know that area of the plate A = 2x3 = 6 m
 
and depth of centre of the plate (average depth) 6+4/2 = 5m 
 
Total thrust on the plate F= pressure*area = hpg*6= 10^3x9.8x6x5 = 294x10'N